OpenStudy (anonymous):

How many different kinds of punch can be made from 2 of the following: cranberry juice, apple juice, orange juice, and passion fruit juice? it be 4!x2!/2 right?

OpenStudy (amistre64):

4 choose 2 i think

OpenStudy (amistre64):

4 pick 2 would have doubles in it

OpenStudy (anonymous):

what do you mean?

OpenStudy (amistre64):

i mean that 4 P 2 = 12 and that 4 C 2 = 6

OpenStudy (amistre64):

4 P 2 counts apple + cran separate from cran + apple

OpenStudy (amistre64):

4 C 2 weeds out the duplicity

OpenStudy (anonymous):

don't think they have that option

OpenStudy (amistre64):

\[4P2=\frac{4!}{(4-2)!}\] \[4C2=\frac{4!}{2!(4-2)!}\]

OpenStudy (anonymous):

is that minus sign suppose to be a multiplication sign?

OpenStudy (amistre64):

no, its spose to be a minus sign; another way to write it is to simplify it as: \(\cfrac{4!}{2!2!}\)

OpenStudy (anonymous):

how did you simplify it?

OpenStudy (amistre64):

(4-2)! = 2!

OpenStudy (anonymous):

ok thanks ima try next one by myself now

OpenStudy (amistre64):

good luck